OpenStudy (anonymous):

4. Express the series below in summation notation for the specified number of terms. a. 2 + 4 + 8 + 16

OpenStudy (anonymous):

\[\sum_{n=1}^{4} 2(n-1)\]

OpenStudy (anonymous):

@ganeshie8 would this equation be like this

ganeshie8 (ganeshie8):

close, but not right

ganeshie8 (ganeshie8):

plugin n = 1, do u get first term ?

ganeshie8 (ganeshie8):

2(1-1) 2(0) 0 NOPE

ganeshie8 (ganeshie8):

check again

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[\sum_{n=2}^{4} 2(n-1)\]

ganeshie8 (ganeshie8):

not exactly, n = 2 to 4 is just three terms, but u have four terms in ur given series right ?

ganeshie8 (ganeshie8):

wait a sec, the given sequence is GEOMETRIC, right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

find out the \(n^{th}\) term of series first

ganeshie8 (ganeshie8):

2 + 4 + 8 + 16 first term, \(a = 2\) common ratio, \(r = 2\) so, the \(n^{th}\) term is ?

OpenStudy (anonymous):

2

ganeshie8 (ganeshie8):

dont remember the \(n^{th}\) term for geometric series ?

ganeshie8 (ganeshie8):

\(\large a_n = ar^{n-1}\)

ganeshie8 (ganeshie8):

plugin the values

OpenStudy (anonymous):

2(2)^n-1

ganeshie8 (ganeshie8):

yes, so the sum notation can be written as : \(\large \sum \limits_{n=1}^4 2(2)^{n-1}\)

ganeshie8 (ganeshie8):

simplifying gives u : \(\large \sum \limits_{n=1}^4 2^{n}\)

OpenStudy (anonymous):

could we write the equation either way

ganeshie8 (ganeshie8):

either way is fine